# Unifying Feature-Based Explanations with Functional ANOVA and Cooperative Game Theory

### v1

@@ F.3 Experimental Setup
We compute explanations for the models and datasets described in Sections F.2 and F.1 based on the shapiq (Muschalik et al., 2024a, ) python library. To do so, we create three kinds of cooperative explanation games, a local explanation game (see Definition 4), a global risk game (see Definition 6), and a sensitivity explanation game (see Definition 5). Both the global risk and global sensitivity explanation games are based on a collection of independent local explanation games. The local explanation game consists of a model, a local data point, and an imputation strategy. The imputation strategies follow the definitions of $F_{S}$ in the b-fANOVA, m-fANOVA, and c-fANOVA. To model b-fANOVA, we compute mean feature values from a background dataset for each feature independently and impute missing features with these baseline values. For m-fANOVA, we sample replacement data points directly from a background dataset according to the joint marginal distribution. For the synthetic datasets and known functional dependencies (see Section F.1), we compute the c-fANOVA with the ground-truth conditional distributions. For real-world datasets and c-fANOVA, we rely on the conditional imputer implemented in the shapiq package (Muschalik et al., 2024b, ). For marginal and conditional imputation we average over $512$ sampled replacement data points for each game evaluation. All imputation strategies operate on the test datasets.
@@ G.1 Dissecting the Linear Models on Synthetic Data
Figure 6 shows the local effects for $F_{\text{add}}$. Comparable to $F_{\text{lin}}$, b-fANOVA is neither influenced by the feature distribution ($\rho$) nor by the higher-order interactions. While m-fANOVA is equivalent to b-fANOVA for $F_{\text{lin}}$, it differs for the still additive relationship, when effects are non-linear: The effect of $x_{2}$ for the instance $x=(1,1,1,1)$ becomes 00 since $f_{1}(x)=\mathbb{E}[F_{\text{add}}(x_{2},X_{-2})-\mathbb{E}[F(X)]=x_{2}+\mathbb{E}[X_{1}+X_{3}^{3}]-\mathbb{E}[X_{1}+X_{2}^{2}+X_{3}^{3}]=0$. However, as for b-fANOVA, pure, partial, and full effects are the same and thus indicating that no feature interactions have been learned by the model. Due to accounting for cross-correlation effects, $x_{4}$ shows again an influence for pure and partial effects, when c-fANOVA is used. Moreover, full effects with c-fANOVA show a similar vanishing trend as for $F_{\text{lin}}$ and $F_{\text{int}}$.
@@ G.2 Additional Explanations on Real-World Data
This section describes additional explanation results on real-world benchmark data as shown in Figures 7, 8, 9 and 10.
@@
Figure 7 shows a series of local explanations with pure (a), partial (b), and full (c) higher-order interaction influence for the Titanic dataset. The local instance under analysis was a 49 year old man who embarked from Cherbourg, holding an upper-class ticket with a relatively high fare price. The XGBoost model correctly predicts that he survived the disaster. The explanations indicate that his Sex (male) and Age (49 years) decrease his chances of survival. In contrast, his high fare price (Fa), the port of embarkation (Em), and his upper-class ticket substanitally enhance his survival likelihood. This finding aligns with descriptive analyses of the Titanic dataset, which reveal that passengers who embarked from Cherbourg had the highest survival rate of $0.56$ in comparison to Queenstown with $0.38$ and $0.33$ for Southampton. Notably, the model also highlights increased survival odds due to interactions between Em and Fa, as well as between Em and Age. Both of these second-order interactions are observed across all three higher-order interaction influences (a-c) depicted in Figure 7.
@@
The local explanation game for the IMDB movie review dataset is based on Sundararajan et al., (2020), Tsai et al., (2023), and Muschalik et al., 2024b and we rely on b-fANOVA with the MASK token as baseline. The sentiment scores are within $\left[-1,1\right]$ where values close to $-1$ denote a negative predicted sentiment and values close to $1$ correspond to a positive predicted sentiment. Sentiment scores close to 00 are considered neutral. The target sentence “The acting was bad, but the movie enjoyable” receives a positive predicted sentiment of approximately $0.8117$, and the baseline prediction is also rather positive around $0.5361$. Figure 11 depicts pure (left), partial (middle) and full (right) individual effects (upper row) and interactions up to the second order (lower row) with 2-SVs (2-SII) for the partial effect. All three effects show strong differences, which indicates the presence of higher-order interactions. The pure individual effects highlight the positive influence of “enjoyable”, whereas “bad” and “but” obtain negative influences. Interestingly, “acting” also receives negative influence, which possibly indicates a bias in the prediction. The partial individual effect (SV) shows a stronger positive influence for ”enjoyable”, and a weaker negative influence for “acting” and “bad”. In contrast, “but” now receives a positive partial influence, which indicates its involvement in higher-order interactions. In fact, “but” and “enjoyable” receives a strong positive interaction, as well as “bad” and “but”. Lastly, full effects show predominantly positive influences, which indicates the presence of many positive higher-order interactions.
@@
Figure 7: Local Explanation Game: Pure (a), partial (b), and full (c) feature influences for an XGBoost model fitted on Bike sharing. The first row depicts the individual explanations where (b) corresponds to interventional SHAP. The second row shows the pure, partial (2-SVs, 2-SII) and full interactions up to the second order. The local instance, which is explained has the following feature values: : $x_{\text{Sex}}=\text{male}$, $x_{\text{Age}}=49$, $x_{\text{SibSp}}=1$, $x_{\text{Parch}}=0$, $x_{\text{Fare}}=56.9292$, $x_{\text{Embarked}}=\text{C}$, $x_{\text{Title}}=\text{Rare Title}$, and $x_{\text{FsizeD}}=\text{small}$.
@@
Figure 8: Global Risk Game: Pure (a), partial (b), and full (c) feature influences for a neural network fitted on California housing. The first row depicts the individual explanations where (b) corresponds to SAGE and (c) corresponds to PFI. The second row shows the pure, partial (2-SVs,2-SII) and full interactions up to the second-order. Importance is defined according to the global risk game by the change in performance measured by the mean-squared error. This global risk game is based on 500 randomly drawn local explanation games from the test dataset.
@@
Figure 9: Global Risk Game: Pure (a), partial (b), and full (c) feature influences for a neural network fitted on Bike sharing. The first row depicts the individual explanations where (b) corresponds to SAGE and (c) corresponds to PFI. The second row shows the pure, partial (2-SVs, 2-SII) and full interactions up to the second-order. Importance is defined according to the global risk game by the change in performance measured by the mean-squared error. This global risk game is based on 100 randomly drawn local explanation games from the test dataset.
Pure, Order 1
Refer to captionRefer to captionRefer to caption
(a) Pure
S
Refer to captionRefer to captionRefer to caption
(b) Partial
Superset Measure $\Upsilon^{2}$
Refer to captionRefer to captionRefer to caption
(c) Full
Figure 10: Global Sensitivity Game: Pure (a), partial (b), and full (c) feature influences for an XGBoost model fitted on Bike sharing. The first row depicts the individual explanations, the second row interactions, and the third row the joint effects up to order $2$. The column (c) corresponds to the superset measure $\Upsilon^{2}$. This global risk game is based on 500 randomly drawn local explanation games from the test dataset.
@@
Figure 11: Local Explanation Game: Pure (a), partial (b), and full (c) feature influences for a sentiment analysis language model fine-tuned on the IMDB dataset (Tsai et al.,, 2023). The first row depicts the individual explanations where (b) corresponds to baseline SHAP using the MASK token, i.e. b-fANOVA. The second row shows pure, partial (2-SVs, 2-SII) and full interactions up to the second-order. The sentence The acting was bad but the movie enjoyable receives a rather positive sentiment score of $0.8117$. Note that the model predicts a score of $0.5361$ for the empty set.

### v2

@@ F.3 Experimental Setup
We compute explanations for the models and datasets described in Sections F.2 and F.1 based on the shapiq (Muschalik et al., 2024a, ) python library. To do so, we create three kinds of cooperative explanation games, a local explanation game (see Definition 4), a global risk game (see Definition 6), and a global sensitivity game (see Definition 5). Both the global risk and global sensitivity games are based on a collection of independent local explanation games. The local explanation game consists of a model, a data point, and an imputation strategy. The imputation strategies follow the definitions of $F_{S}$ in the b-fANOVA, m-fANOVA, and c-fANOVA. To model b-fANOVA, we compute mean feature values from a background dataset for each feature independently and impute missing features with these baseline values. For m-fANOVA, we sample replacement data points directly from a background dataset according to the joint marginal distribution. For the synthetic datasets and known functional dependencies (see Section F.1), we compute the c-fANOVA with the ground-truth conditional distributions. For real-world datasets and c-fANOVA, we rely on the conditional imputer implemented in the shapiq package (Muschalik et al., 2024b, ). For marginal and conditional imputation we average over $512$ sampled replacement data points for each game evaluation. All imputation strategies operate on the test datasets.
@@ G.1 Dissecting the Linear Models on Synthetic Data
Figure 6 shows the local effects for $F_{\text{add}}$. Comparable to $F_{\text{lin}}$, b-fANOVA is neither influenced by the feature distribution ($\rho$) nor by the higher-order interactions. While m-fANOVA is equivalent to b-fANOVA for $F_{\text{lin}}$, it differs for the still additive relationship, when effects are non-linear: The effect of $x_{2}$ for the instance $x=(1,1,1,1)$ becomes 00 since $f_{1}(x)=\mathbb{E}[F_{\text{add}}(x_{2},X_{-2})]-\mathbb{E}[F(X)]=x_{2}+\mathbb{E}[X_{1}+X_{3}^{3}]-\mathbb{E}[X_{1}+X_{2}^{2}+X_{3}^{3}]=0$. However, as for b-fANOVA, pure, partial, and full effects are the same and thus indicating that no feature interactions have been learned by the model. Due to accounting for cross-correlation effects, $x_{4}$ shows again an influence for pure and partial effects, when c-fANOVA is used. Moreover, full effects with c-fANOVA show a similar vanishing trend as for $F_{\text{lin}}$ and $F_{\text{int}}$.
@@ G.2 Additional Explanations on Real-World Data
This section describes additional explanation results on real-world benchmark data as shown in Figures 7, 9, 10 and 8.
@@
Figure 7 shows a series of local explanations with pure (a), partial (b), and full (c) higher-order interaction influence for the Titanic dataset. The local instance under analysis was a 49 year old man who embarked from Cherbourg, holding an upper-class ticket with a relatively high fare price. The XGBoost model correctly predicts that he survived the disaster. The explanations indicate that his Sex (male) and Age (49 years) decrease his chances of survival. In contrast, his high fare price (Fa), the port of embarkation (Em), and his upper-class ticket substantially enhance his survival probability. This finding aligns with descriptive analyses of the Titanic dataset, which reveal that passengers who embarked from Cherbourg had the highest survival rate of $0.56$ in comparison to Queenstown with $0.38$ and $0.33$ for Southampton. Notably, the model also highlights increased survival odds due to interactions between Em and Fa, as well as between Em and Age. Both of these second-order interactions are observed across all three higher-order interaction influences (a-c) depicted in Figure 7.
Titanic Data - Global Sensitivity Game.
@@
The local explanation game for the IMDB movie review dataset is based on Sundararajan et al., (2020), Tsai et al., (2023), and Muschalik et al., 2024b and we rely on b-fANOVA with the MASK token as baseline. The sentiment scores are within $\left[-1,1\right]$ where values close to $-1$ denote a negative predicted sentiment and values close to $1$ correspond to a positive predicted sentiment. Sentiment scores close to 00 are considered neutral. The target sentence “The acting was bad, but the movie enjoyable” receives a positive predicted sentiment of approximately $0.8117$, and the baseline prediction is also rather positive around $0.5361$. Figure 11 depicts pure (left), partial (middle), and full (right) individual effects (upper row) and interactions up to the second order (lower row) with 2-SVs (2-SII) for the partial effect. All three effects show strong differences, which indicates the presence of higher-order interactions. The pure individual effects highlight the positive influence of “enjoyable”, whereas “bad” and “but” obtain negative influences. Interestingly, “acting” also receives negative influence, which possibly indicates a bias in the prediction. The partial individual effect (SV) shows a stronger positive influence for ”enjoyable”, and a weaker negative influence for “acting” and “bad”. In contrast, “but” now receives a positive partial influence, which indicates its involvement in higher-order interactions. In fact, “but” and “enjoyable” receives a strong positive interaction, as well as “bad” and “but”. Lastly, full effects show predominantly positive influences, which indicates the presence of many positive higher-order interactions.
@@
Figure 7: Local Explanation Game: Pure (a), partial (b), and full (c) feature influences for an XGBoost model fitted on Titanic. The first row depicts the individual explanations where (b) corresponds to interventional SHAP. The second row shows the pure, partial (2-SVs, 2-SII) and full interactions up to the second order. The local instance, which is explained has the following feature values: : $x_{\text{Sex}}=\text{male}$, $x_{\text{Age}}=49$, $x_{\text{SibSp}}=1$, $x_{\text{Parch}}=0$, $x_{\text{Fare}}=56.9292$, $x_{\text{Embarked}}=\text{C}$, $x_{\text{Title}}=\text{Rare Title}$, and $x_{\text{FsizeD}}=\text{small}$.
Pure, Order 1
Refer to captionRefer to captionRefer to caption
(a) Pure
S
Refer to captionRefer to captionRefer to caption
(b) Partial
Superset Measure $\Upsilon^{2}$
Refer to captionRefer to captionRefer to caption
(c) Full
Figure 8: Global Sensitivity Game: Pure (a), partial (b), and full (c) feature influences for an XGBoost model fitted on Titanic. The first row depicts the individual explanations, the second row interactions, and the third row the joint effects up to order $2$. The column (c) corresponds to the superset measure $\Upsilon^{2}$. This global risk game is based on 500 randomly drawn local explanation games from the test dataset.
@@
Figure 9: Global Risk Game: Pure (a), partial (b), and full (c) feature influences for a neural network fitted on California housing. The first row depicts the individual explanations where (b) corresponds to SAGE and (c) corresponds to PFI. The second row shows the pure, partial (2-SVs,2-SII) and full interactions up to the second-order. Importance is defined according to the global risk game by the change in performance measured by the mean squared error. This global risk game is based on 500 randomly drawn local explanation games from the test dataset.
@@
Figure 10: Global Risk Game: Pure (a), partial (b), and full (c) feature influences for a neural network fitted on Bike sharing. The first row depicts the individual explanations where (b) corresponds to SAGE and (c) corresponds to PFI. The second row shows the pure, partial (2-SVs, 2-SII) and full interactions up to the second-order. Importance is defined according to the global risk game by the change in performance measured by the mean squared error. This global risk game is based on 100 randomly drawn local explanation games from the test dataset.
@@
Figure 11: Local Explanation Game: Pure (a), partial (b), and full (c) feature influences for a sentiment analysis language model fine-tuned on the IMDB dataset (Tsai et al.,, 2023). The first row depicts the individual explanations where (b) corresponds to baseline SHAP using the MASK token, i.e., b-fANOVA. The second row shows pure, partial (2-SVs, 2-SII) and full interactions up to the second-order. The sentence The acting was bad, but the movie enjoyable receives a rather positive sentiment score of $0.8117$. Note that the model predicts a score of $0.5361$ for the empty set.